BlueWolf08:

When solving negative one over eight (x + 35) = −7, what is the correct sequence of operations? Multiply each side by negative one over eight , add 35 to each side Multiply each side by negative one over eight , subtract 35 from each side Multiply each side by −8, subtract 35 from each side Multiply each side by −8, add 35 to each side Homework Help Meh Plz I got 3 more after this one sooo... AHHHHHHHHHH HALPPPPP IS IT C?!

BlueWolf08:

Why dis Gotta be so dang hard!?!?

ellapayne:

Is it subtracting to get the answer that you need?

ellapayne:

if so, then that is your answer choice.

BlueWolf08:

Uhh

BlueWolf08:

Ok Well just so ya know im in 7th grade so its not like easy

ellapayne:

oh

BlueWolf08:

oop

jhonyy9:

\[\frac{ -1 }{ 8 }\left( x +35 \right) = -7\]

BlueWolf08:

which answer choice is that?

BlueWolf08:

C.?

jhonyy9:

what you need to do to eliminate the denominator ?

jhonyy9:

yes of cors and the minus from both sides ?

BlueWolf08:

yeah this is an example... First method: treat (x − 4) as a quantity 2(x−4) = −92 2(x−4) 2 = −92 2 2 2 (x−4) = −92 2 x−4 = −46 x−4 +4 = −46 +4 x = −42 Second method: simplify using the distributive property 2(x−4) = −92 2x−8 = −92 2x−8 +8 = −92 +8 2x = −84 2x 2 = −84 2 x = −42

jhonyy9:

you wrote above choice C. - sure ?

jhonyy9:

in this way what will get ?

BlueWolf08:

uhmm

BlueWolf08:

I-

jhonyy9:

multiplie both sides by -8 and so will get x+35 = 56

jhonyy9:

yes ?

BlueWolf08:

Yeah

jhonyy9:

and now for x

jhonyy9:

what say choice C. ?

BlueWolf08:

would be... C because you divide not multiply

jhonyy9:

yes choice C but you need calcule the x value

jhonyy9:

x = ?

BlueWolf08:

umm

BlueWolf08:

idk

jhonyy9:

x +35 = 56 x = ?

BlueWolf08:

something negative I guess and a fraction

BlueWolf08:

wait so hold on

BlueWolf08:

I think I got it

jhonyy9:

x = ?

jhonyy9:

x+35 = 56 subtract from both sides 35 so what will get ?

BlueWolf08:

21

BlueWolf08:

?????

BlueWolf08:

Is it C

BlueWolf08:

umm.. sooo C

jhonyy9:

yes C

BlueWolf08:

YAY!!!

BlueWolf08:

Lmaooo

jhonyy9:

bye

BlueWolf08:

Byee